# ChessGPT: Bridging Policy Learning and Language Modeling

### v1

@@ D.1.3 Compute resources
D.2 Evaluation details
D.2.1 Evaluation on chess modeling ability
@@
D.2.2 Evaluation on task construction
@@
Appendix E Qualitative Analysis
@@ E.1 Experimental set-up
E.2 General knowledge
In this section, we ask each model some questions about general knowledge of chess. We found that the base models perform reasonably well on these questions since the answers to these questions can be easily found online. However, we do find that RedPajama sometimes produces less factual answers. Detailed examples can be found in tables 17, 19 and 16.
| {tcolorbox}[title = Prompt, fonttitle = , fontupper = , fontlower = ] The quickest possible checkmate is in: (Answer: 2 moves, Fool’s mate) |
{bbox}ChessGPT-Base
{gbox}ChessGPT-Chat
{rbox}RedPajama
| {tcolorbox}[title = Prompt, fonttitle = , fontupper = , fontlower = ] What is it called when a player can’t defend an attack against their king? (Answer: checkmate) |
{bbox}ChessGPT-Base
{gbox}ChessGPT-Chat
{rbox}RedPajama
| {promptbox} What determines which player starts first in a chess game? (Answer: white) |
{bbox}ChessGPT-Base
{gbox}ChessGPT-Chat
{rbox}RedPajama
| {tcolorbox}[title = Prompt, fonttitle = , fontupper = , fontlower = ] Does DeepBlue defeat Garry Kasparov? (Answer: yes) |
{bbox}ChessGPT-Base
{gbox}ChessGPT-Chat
{rbox}RedPajama
E.3 Chess modeling
@@
| {tcolorbox}[title = Prompt, fonttitle = , fontupper = , fontlower = ] 1.e4 e5 2.f4, What is the name of this opening? (Answer: King’s Gambit) |
{bbox}ChessGPT-Base
{gbox}ChessGPT-Chat
{rbox}RedPajama
@@
| {tcolorbox}[title = Prompt, fonttitle = , fontupper = , fontlower = ] 1. e4 e5 2. Nf3 d6 3. d4 exd4 4. Nxd4 Nf6 5. Nc3 Qe7 6. Bd3 d5 7. O-O dxe4 8. Re1 Be6 9. Nxe6 fxe6 10. Bxe4 Nxe4 11. Nxe4 Nd7 12. Bg5 Qb4 13. Qg4 Qd4 14. Qxe6+ Be7 15. Convert the PGN to FEN (Answer: r3k2r/pppnb1pp/4Q3/6B1/3qN3/8/PPP2PPP/R3R1K1 w kq - 1 15) |
{bbox}ChessGPT-Base
{gbox}ChessGPT-Chat
{rbox}RedPajama
@@
| {tcolorbox}[title = Prompt, fonttitle = , fontupper = , fontlower = ] Convert the following UCI to FEN: g2g3 d7d6 f1g2 e7e5 g1f3 g8f6 d2d4 b8 (Answer: rnbqkb1r/ppp2ppp/3p1n2/4p3/3P4/5NP1/PPP1PPBP/RNBQK2R b KQkq - 0 4) |
{bbox}ChessGPT-Base
{gbox}ChessGPT-Chat
{rbox}RedPajama
@@
| {promptbox} 1. e4 e5 2. Nf3 d6 3. d4 exd4 4. Nxd4 Nf6 5. Nc3 Qe7 6. Bd3 d5 7. O-O dxe4 8. Re1 Be6 9. Nxe6 fxe6 10. Bxe4 Nxe4 11. Nxe4 Nd7 12. Bg5 Qb4 13. Qg4 Qd4 14. Qxe6+ Be7 15. Generate a step-by-step analysis |
{bbox}ChessGPT-Base
{gbox}ChessGPT-Chat
{rbox}RedPajama
Appendix F Potential directions
@@
Dataset Augmentation and Fine-tuning: Researchers can explore dataset augmentation techniques to further enhance the diversity and size of the provided dataset. By introducing variations in game conditions, player strategies, or opening positions, researchers can create augmented datasets that can be used to fine-tune existing language models or train new models specifically tailored for chess-related tasks. This can potentially lead to improved performance in state tracking, value judgment, and policy evaluation..
@@
Transfer Learning to Chess Variants: The dataset and benchmark provided in this study can serve as a valuable resource for transfer learning experiments to chess variants. Researchers can leverage the knowledge learned from the base chess task and apply it to chess variants such as Fischer Random Chess or Chess960. By fine-tuning or adapting pre-trained language models on the provided dataset, researchers can explore the capabilities of these models in handling different chess variants, including state tracking, value judgment, and policy evaluation
@@
Multi-modal Approaches: Researchers can explore multi-modal approaches that combine textual and visual information for chess-related tasks. By incorporating board visualizations, game position images, or move sequence visualizations along with textual inputs, researchers can develop models that leverage both textual and visual cues to improve state tracking, value judgment, and policy evaluation in chess. This can open up avenues for multi-modal analysis and understanding of chess games, allowing models to capture and reason over both textual and visual representations simultaneously.
@@
Chess Education and Tutorial Systems: The dataset can be utilized to develop educational tools and tutorial systems for chess players of different skill levels. Researchers can leverage the language model’s expertise in state tracking, value judgment, and policy evaluation to provide interactive and personalized learning experiences. By tailoring the tutorial content and feedback based on individual player performance, researchers can create intelligent systems that assist in skill development and strategic improvement in chess.
@@
Appendix G Limitations and Potential Societal Impact

### v2

@@ F.1.3 Compute resources
Appendix G Evaluation details
G.1 Evaluation task examples and plots
Here we show task examples and plots for our evaluation tasks. Basically, the evaluation tasks consist of three parts: Task Prefix, which can be regarded as a description of the task and is also the main prompt we use for LLMs. Input, which is the question and the input of LLMs. And Target, which contains the answer of the question for exact string match tasks, or target score, which provides the score for each available answer for multi-choice tasks.
1. e4 e6 2. Ke2 d5 3. e5 c5 4. f4 Nc6 5. Nf3 Qb6 6. g4 Bd7 7. h4 Nge7 8. c3 Ng6 9. d4 cxd4 10. cxd4 Be7 11. Kf2 O-O 12. h5 Nh8 -----> Bxh7+ 13. Be3 Qxb2+ 14. Kg3 Qxa1 15. Bd3 Qxa2 16. Rh2 Qa1 17. Qc2 Nb4 18.
G.1.1 Evaluation on chess modeling ability
@@
G.1.2 Evaluation on task construction
@@
Appendix H Qualitative Analysis
@@ H.1 Experimental set-up
H.2 General knowledge
In this section, we ask each model some questions about general knowledge of chess. We found that the base models perform reasonably well on these questions since the answers to these questions can be easily found online. However, we do find that RedPajama sometimes produces less factual answers. Detailed examples can be found in tables 19, 21 and 18.
H.3 Chess modeling
@@
Appendix I ChessCLIP visualization
We present a visualization demo to better illustrate ChessCLIP’s capability. Here we choose the chess opening as a test scenario and choose ten different chess openings. Ten chess opening PGNs and their corresponding names are arranged in sequence, so the diagonal cells in the entire similarity matrix should have the highest similarity. We present such a similarity matrix generated by ChessCLIP in Figure 6. The results exactly correspond to our expectations, which successfully validate the effectiveness of our ChessCLIP model.
Appendix J Potential directions
@@ Appendix I ChessCLIP visualization
Dataset Augmentation and Fine-tuning: Researchers can explore dataset augmentation techniques to further enhance the diversity and size of the provided dataset. By introducing variations in game conditions, player strategies, or opening positions, researchers can create augmented datasets that can be used to fine-tune existing language models or train new models specifically tailored for chess-related tasks. This can potentially lead to improved performance in state tracking, value judgement, and policy evaluation.
@@
Transfer Learning to Chess Variants: The dataset and benchmark provided in this study can serve as a valuable resource for transfer learning experiments to chess variants. Researchers can leverage the knowledge learned from the base chess task and apply it to chess variants such as Fischer Random Chess or Chess960. By fine-tuning or adapting pre-trained language models on the provided dataset, researchers can explore the capabilities of these models in handling different chess variants, including state tracking, value judgement, and policy evaluation
@@
Multi-modal Approaches: Researchers can explore multi-modal approaches that combine textual and visual information for chess-related tasks. By incorporating board visualizations, game position images, or move sequence visualizations along with textual inputs, researchers can develop models that leverage both textual and visual cues to improve state tracking, value judgement, and policy evaluation in chess. This can open up avenues for multi-modal analysis and understanding of chess games, allowing models to capture and reason over both textual and visual representations simultaneously.
@@
Chess Education and Tutorial Systems: The dataset can be utilized to develop educational tools and tutorial systems for chess players of different skill levels. Researchers can leverage the language model’s expertise in state tracking, value judgement, and policy evaluation to provide interactive and personalized learning experiences. By tailoring the tutorial content and feedback based on individual player performance, researchers can create intelligent systems that assist in skill development and strategic improvement in chess.
@@
Appendix K Limitations and Potential Societal Impact